I am really having trouble transferiing this question from words to the actual problem please help,
Find 2 consecutive odd integers such that twice the greater is 17 more than the lesser.

Answer by stanbon(75887)

Find 2 consecutive odd integers such that twice the greater is 17 more than the lesser.
---------------------
smaller: 2x-1
larger: 2x+1
-------------------------
Equation:
2(2x+1) = 2x-1 + 17
4x+2 = 2x + 16
2x = 14
x = 7
----------------
smaller: 2x-1 = 2*7-1 = 13
lafter: 2x+1 = 2*7_1 = 15
===============================
Cheers,
Stan H.